Deedee invested $2000, part at 7% per annum and the rest at 8% per annum. After one year, the interest earned on the 7% investment was $50 more than the interest on the 8% investment. How much (money) did she invest a each rate?

Answer by richwmiller(17219)

x+y=2000,
.07x=.08y+50,
x+y=2000
.07(2000-y)-.08y=50
investment
x=1400 at 7%, y=600 at 8%
interest earned
98-48=50
$146 total interest